### v4

# Successor-Generator Planning with LLM-generated Heuristics

Abstract
Heuristics are a central component of deterministic planning, particularly in domain-independent settings where general applicability is prioritized over task-specific tuning. This work revisits that paradigm in light of recent advances in large language models (LLMs), which enable the automatic synthesis of heuristics directly from problem definitions – bypassing the need for handcrafted domain knowledge. We present a method that employs LLMs to generate problem-specific heuristic functions from planning tasks specified through successor generators, goal tests, and initial states written in a general-purpose programming language. These heuristics are compiled and integrated into standard heuristic search algorithms, such as greedy best-first search. Our approach achieves competitive, and in many cases state-of-the-art, performance across a broad range of established planning benchmarks. Moreover, it enables the solution of problems that are difficult to express in traditional formalisms, including those with complex numeric constraints or custom transition dynamics. We provide an extensive empirical evaluation that characterizes the strengths and limitations of the approach across diverse planning settings, demonstrating its effectiveness.
Introduction
Deterministic AI planning involves finding a sequence of actions that moves an agent from an initial state to a goal state, given a formal description of the environment’s dynamics (ghallab-et-al:2004). Because state spaces are usually exponential or even infinite, effective search strategies require heuristics – functions that estimate the distance to the goal and help prioritize promising states (pearl:1984; bonet2001planning). Heuristic search has therefore become a central paradigm in classical planning, underpinning many of the most successful planners developed over the past decades (hoffman:jair-2003; helmert:jair-2006; scala-et-al:ijcai-2016; aldinger-nebel:socs-2017).
Classical planning has traditionally focused on domain-independent heuristics, which offer broad applicability and avoid the need for task-specific design (bonet2001planning). However, the performance of such heuristics can vary with the representation language and the structural characteristics of the domain. In more expressive settings, general heuristics may encounter challenges and need adaptation to remain effective across different versions of pddl (mcdermott-et-al:tr-1998; fox-long:jair-2003; edelkamp-hoffmann:tr-2004) (e.g., cf. classical vs. numeric settings (hoffman:aimag-2001; hoffman:jair-2003; helmert-domshlak:icaps-2009; kuroiwa-et-al:jair-2022)). For further discussion of current limitations within pddl see edelkamp:icapsws-2003; rintanen:aaai-20215.
In this work, we propose a method for automatically generating problem-specific heuristic functions from formal problem definitions using LLMs. The generated heuristics are then used by a sound search algorithm, such as greedy best-first search (GBFS). This setup preserves the systematic and provably correct behavior of classical search methods, while making heuristic construction fully automatic and eliminating the need for expert-designed heuristics.
We represent each planning task using three components, implemented in a general-purpose programming language: a successor generator, a goal test, and an initial state (russell-norvig:1995; guan-etal:nips-2023; oswald-etal:icaps-2024). We refer to this representation as an Explicit Successor Generator (ESG). These ESG components, together with a prompt, are provided as input to the LLM, which returns a heuristic function (expressed in the same programming language) tailored to the structure of the given problem. The heuristic is then compiled and used to guide search without any further interaction with the LLM.
This approach offers several benefits. It removes the need for repeated LLM calls during planning—a source of significant inefficiency in prior work (tos:nips-2024; valmeekam-etal:corr-2022; kambhampati-etal:icml-2024). It also provides transparent search components: both the LLM-generated heuristic and the task specification are expressed in a general-purpose programming language (Rust in this work), allowing inspection, debugging, and testing. In addition, it accommodates planning tasks that involve constructs difficult to model in traditional frameworks, including recursive goals, intricate numeric conditions, the creation of new variables, or custom transition logic.
We show that LLM-generated heuristic can strengthen general-purpose ESG planning. Our evaluation includes a set of standard numerical planning benchmarks represented in the ESG form (using Rust), as well as new benchmarks that are challenging to encode in pddl. Across many domains, the resulting heuristics achieve strong performance, matching the level of state-of-the-art numeric planners in numerous cases. Furthermore, the method is able to solve planning tasks that currently lack formal encodings in existing planning languages.
Related Work
This paper revisits AI planning by heuristic search, by replacing domain-independent heuristics with LLM-generated domain-specific ones. Below, we review related work on LLMs and planning. For a more comprehensive overview of LLMs as planning modelers, see the surveys by (tantakoun-etal:corr-2025) and by (aghzal2025surveylargelanguagemodels).
Planning with LLMs
(valmeekam-etal:corr-2022) provides an important early evaluation of LLM-based planning, demonstrating that even on simple benchmarks, LLMs do not consistently exhibit reliable planning behavior. While it seems that LLMs cannot yet generate plans efficiently, our work shows that they can be used effectively as a component of a larger planning paradigm. Namely, LLMs can generate heuristics and search components rather than perform planning directly.
A closely related approach is presented by (tos:nips-2024), who argue that existing LLM-based planning methods are neither sound nor complete and incur high computational costs due to repeated LLM calls. They propose an alternative, Thought of Search (ToS), in which LLMs are used to generate symbolic search components—successor functions and goal tests. These can then be executed efficiently without further model queries. Although their work focuses on small, finite search problems such as the 24 game and Mini Crosswords, they provide compelling evidence that such an approach can drastically reduce computational cost and improve correctness (solved by BFS). The work shows that with respect to planning their approach overperforms the planning directly with LLM approaches such as: Chain-of-Thought (wei2022chain), ReAct (react:iclr-2023), ReWOO (rewoo:corr-2023), Reasoning via Planning (rap:emnkp-2023), Tree of Thoughts (tot:nips-2023), Graph of Thoughts (got:aaai-2024), Reflexion (reflection:nips-2023), Algorithm of Thought (aot:icml-2024). The ToS approach runs a breadth-first search (BFS) using components—successor functions and goal tests generated by LLM. Our approach extends this intuition by employing heuristic search methods to address problems of significantly greater complexity than those solvable through naive BFS—specifically in AI planning, where transition systems often become infinite due to the presence of unbounded numeric fluents. Thus, generating effective heuristics is particularly crucial, underscoring that the symbolic utilization of LLMs can be successfully applied beyond toy domains.
A recent study in a similar vein to our work is presented in (correa2025classical). While their approach targets planning tasks specified in pddl, ours departs from this framework by operating directly on successor generators and goal-check functions. This shift affords greater representational flexibility, enabling the modeling of tasks that PDDL cannot express effectively. However, abandoning pddl’s structured representation also limits our ability to exploit domain structure for heuristic derivation.
Heuristic Generation Outside of Planning
LLMs have also been explored for heuristic generation beyond automated planning, notably in domains such as online bin packing and TSP (liu2024evolution), as well as other combinatorial optimization problems (ye2024reevo). These approaches typically combine LLMs with evolutionary algorithms to evolve heuristics that guide the optimization process. However, due to fundamental differences in problem structure, these methods are not directly applicable to planning tasks without substantial adaptation.
Problem Generation
Recent work has explored using LLMs to generate domains and problem descriptions using pddl2.1 (fox-long:jair-2003), among such works are (guan-etal:nips-2023; liu-etal:corr-2023; oswald-etal:icaps-2024). For instance, (silver-etal:aaai-2024) demonstrated that GPT-4 can output plans from parsed pddl2.1 inputs using simple, non-search-based strategies. (guan-etal:nips-2023) proposed an approach in which pddl2.1 actions are generated one at a time via LLM queries, refined using human feedback. A related concept appears in (zhou-etal:iclr-2024), who use LLMs to generate, execute, and refine Python code for solving mathematical reasoning problems. Their findings support the idea that LLMs can produce verifiable code with minimal manual intervention. To reduce human involvement even further, we propose generating Rust code instead of Python. Since Rust is a compiled language with strong static typing, much of the burden of code verification can be offloaded to the compiler.
Code Generation Our approach relies on LLMs not to solve problems directly, but to generate search guidance in the form of state evaluation heuristic code. (zhou-etal:iclr-2024) pursued a similar strategy in mathematical domains, showing that LLMs can generate and iteratively improve executable code. Our work builds on this insight, but targets search guidance functions within a planning framework. The viability of this approach is supported by recent advances in program synthesis and LLM-based code generation (madaan-etal:nips-2023; zhang-etal:acl-2023; chen-etal:iclr-2024; muennighoff-etal:iclr-2024; zhong-etal:aclf-2024), which show that language models are increasingly capable of producing correct, type-safe, and semantically meaningful programs.
Deterministic AI Planning
AI planning refers to the problem of computing a sequence of actions that transforms a known initial state into one that satisfies a specified goal condition. Unlike reinforcement learning, which learns behavior through trial-and-error interaction with an environment, AI planning assumes full knowledge of the state transition dynamics and the goal. The planner searches the combinatorial space of states reachable via actions to construct a valid plan.
Formally, a deterministic planning problem can be defined as a tuple $\Pi=\langle V,A,T,s_{0},G\rangle$, where: $V$ is a set of variables with either finite or numeric domains, $A$ is a set of symbolic actions, $T$ is a deterministic transition function, $s_{0}\in S$ is the initial state, and $G$ is the goal description. A state $s\in S$ is a full assignment over the variables in $V$. Since the state space $S$ is typically exponential in $|V|$ (or even infinite in the numeric case), it is not explicitly represented in the input. The goal description $G$ is usually represented as a set of conditions. A state $s_{\ast}\in S$ is considered a goal state if it satisfies $G$, i.e., for each $\varepsilon\in G$ the state $s_{\ast}$ satisfies $\varepsilon$. The transition function $T$ maps a state-action pair to the resulting state, i.e., $T(s,a)=s^{\prime}$. A solution (or plan) for $\Pi$ is a sequence of actions $\langle a_{1},\ldots,a_{n}\rangle$ such that applying them in order leads from $s_{0}$ to a goal state $s_{n}$.
Planning problems can be classified as either satisficing, where the goal is to find any valid plan, or optimal, where the goal is to find a cost-minimizing plan. In this work, we focus on satisficing planning, which prioritizes informative heuristics over theoretically sound ones (e.g., admissibility).
Solutions to planning problems are typically found via heuristic search algorithms such as GBFS or A* (hart-et-al:ieee-ssc-1968; pearl:1984). These algorithms are guided by heuristic functions that estimate the cost or distance from a given state to the goal (pearl:1984). A heuristic function $h:S\to\mathbb{R}_{\geq 0}$ provides such estimates, often derived using domain-independent approximations, such as relaxations of the planning problem. Traditional heuristics are based on symbolic representations and formal action models (e.g., in pddl2.1(fox-long:jair-2003), but recent work also explores learned heuristics (oswald-etal:icaps-2024; chen-etal:aaai-2024). Informative heuristics greatly reduce the number of explored states, making them essential for practical planning. In this work, we use an LLM to generate a heuristic function based on the formal problem description written in a general-purpose programming language (Rust). This heuristic is integrated into a GBFS planner to obtain satisficing solutions.
Problem Input: pddl2.1 Input: General Description Translate pddl to Rust Generate Rust Functions Generate Heuristic Code via LLM Compile into GBFS and Run Verify with val Verify with Custom Procedure Solution
Figure 1: Procedure flowchart. A general problem description is manually written using Rust components—successor generator, goal test, and initial state—which are then fed into the system. For problems already described in pddl2.1, the Rust translation is derived directly from the encoding.
Methodology
We start with the overview of our approach. The process begins by accepting two types of input: structured representations provided in pddl2.1 (fox-long:jair-2003) format and general problem descriptions given informally or in non-standardized formats. The problems specified in pddl2.1 are manually translated into Rust functions that implement a successor generator and a goal-testing function. For general problem descriptions, we encode them into Rust according to their specifications. Moreover, the initial and the goal states are specified separately in a JSON file. The manual translation is necessary because current LLMs are not sufficiently reliable for this task, and existing tools lack the maturity required to perform it effectively. The automatic translations to ESG formalism is left for future work.
With the foundational components in place, we employ an LLM to generate the heuristic function. The LLM is provided with the Rust code for the successor generator and goal-testing function, along with a prompt designed to produce a heuristic function in Rust suited to the problem. The prompt [^1] The exact prompt for both variations is given in the Appendix. has three components:
  1. 1.
    Conditioning the model to be a senior engineer in the GPL used (Anam2025PromptEA) and providing the format the resulting heuristic must follow. This component is spread across both the system prompt and the user message [^2] We include guidance to avoid compilation errors, however the guidance could be far better optimized. This optimization, however, is out of scope for this work.
  2. 2.
    Requesting that the LLM generate a heuristic.
  3. 3.
    Providing the model with the GPL domain implementation (the ESG).
As an additional variation of our method, the user message may also contain a fourth component:
  1. 4.
    The JSON representation of the instance to be solved.
We analyze the impact of instance-specificity in the heuristic generation pipeline in the Ablation Study section.
The heuristic, successor generator, and goal-testing functions are then integrated into a standard GBFS framework, where the search is guided by the generated heuristic. To ensure correctness, all solutions are verified using appropriate mechanisms based on the input type. For problems specified in pddl2.1, we rely on a standard pddl verifier—val (howey-etal:ictai-2004), which cross-checks the solution against the formal specification of the problem. For general problem descriptions, we verify by applying the actions in the plan sequentially, verify the preconditions are met, and that the last state fulfills goal-test. For the visualization of the workflow see Fig. 1. This workflow integrates LLM capabilities with automated code generation and classical search techniques, enabling efficient and flexible problem-solving for a wide range of planning domains. Note that the only element of the workflow that still requires human intervention is the specification of the problem.
Baseline Domain-Independent Planners FirstCompilation (fc) UntilSuccess (US) SelfPortfolio-10 (SP-10)
Rust ENHSP-20 NFD MFF GPT Claude GPT Claude GPT Claude
Domain BFS $h^{\text{md}}$ $h^{\text{md}}$ $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ $P(3h||3n)^{\ddagger}$ $h_{\text{LMC}}$ $h_{\text{FF}}$ 4.1 5.1 Sonnet 4.5 4.1 5.1 Sonnet 4.5 4.1 5.1 Sonnet 4.5
[…]
Table 1: Coverage results of the baseline, domain-independent planners, the best portfolio, and the LLM-generated heuristics. We report our results from GPT-4.1, GPT-5.1, and Claude Sonnet 4.5 with configurations FirstCompilation (fc), UntilSuccess (us), and SelfPortfolio-10 (SP-10). All configuration except BFS and MFF (which uses EHCS) use GBFS. Models marked with are reasoning models. For results on all models and configurations see Appendix. The $P(3h||3n)$ portfolio uniformly divides time among three “best” ENHSP heuristics and three “best” novelty heuristics (chen-thiebaux:socs-2024), where the notion of “best” is derived from performance on the same problem set used in our evaluation, rather than being fixed independently of it.
Empirical Evaluation
The experiments were conducted on a 13th Gen Intel® Core™ i9-13900 processor running at 2.00 GHz, supported by a 64-bit operating system and 32 GB of RAM. To ensure consistency and reliability, each experiment was constrained by a time limit of 10 minutes and a memory usage limit of 8 GB. We test our approach in the context of satisficing numeric planning. We have evaluated our approach on the domains from the Numeric International Planning Competition (IPC) 2023 (taitler-etal:aimag-2024). [^3] https://github.com/ipc2023-numeric Note that even problems with three unbounded numeric have infinite size state spaces and in general are undecidable (gnad-et-al:icaps-2023). As a proof-of-concept, we also include two domains that cannot be easily expressed using pddl:
Twin Prime: The initial state comprises a set of integers and buffers, each holding a single integer. The actions include addition, subtraction, multiplication, and integer division between two registers, with results stored in one register. The goal is to produce a twin prime number exceeding a specified threshold in one of the buffers—a goal that is not easily expressible in pddl.
Deterministic Pacman: The game is defined on a grid containing walls, pellets, ghosts, power-ups, and a single Pacman agent. Pacman moves in one of four cardinal directions per turn, consuming pellets and power-ups upon entry into their cell. The goal is to clear all pellets while avoiding collisions with ghosts that behave in a deterministic manner. Consuming a power-up gives Pacman a temporary ability to banish ghosts and avoid harm from them. This domain cannot be easily expressed in pddl or its extensions due to the complex interactions between Pacman, ghosts, and power-ups, which require temporal dynamics and conditional effects beyond pddl’s representational capabilities.
(a) Total Time (including failed generation + run pairs)
(b) Expanded States
Figure 2: Per-instance comparisons of the Total Time (left) and expanded states (right) between GPT-4.1 with SelfPortfolio-10 and $h^{\text{add}}_{\langle\text{B, QB}\rangle}$. Points below the diagonal favor our approach. On problems both solve they seem to provide similar levels of heuristic guidance. [^5] A portfolio between the two achieves a coverage of 304, however this is a post-hoc optimization and not a fair comparison. We do however consider Domain-Independent+LLM a fruitful direction fo future work.
Figure 3: Per-instance comparison of the Total Time (generation + run) (up) and expanded states (down) between GPT-5.1 with and without setting the reasoning_effort to high. Points below the diagonal favor high reasoning effort. Allowing increased reasoning effort moderately increases heuristic quality, but comes at a heavy expense of time.
We evaluated our approach against the following SOTA methods on the IPC domains. All of the planners below support linear expressions in both conditions and assignment effects. As a baseline, we implement a “blind” BFS and GBFS with a manually implemented $h^{\text{md}}$ heuristic (chen-thiebaux:socs-2024), running on a Rust implementation of the problems based on (Domains_Project).
The $h^{\text{md}}$ heuristic estimates the total distance from a state $s$ to satisfying all goal conditions, inspired by the $h^{\text{md}}$ norm:
$h^{\text{md}}(s):=\sum_{\varepsilon\in G}d(\varepsilon,s),$
where $G$ is the set of goal conditions, and $d(\varepsilon,s)$ measures the distance from $s$ to satisfying $\varepsilon$.
[…]
$d(\varepsilon,s):=\inf_{y:\psi(y)\,\bowtie\,0}|\psi(y)-\psi(s)|,$
i.e., the distance from $\psi(s)$ to the nearest satisfying value of $\psi$. chen-thiebaux:socs-2024 proposed this approach for linear constraints in pddl2.1, but this heuristic extends naturally to non-linear formulas, though computing $d$ becomes more expensive.
ENHSP-20 is a Java-based planner whose satisficing configurations use GBFS by (scala-et-al:jair-2020), using the modified version released by chen-thiebaux:socs-2024.666github.com/hstairs/jpddlplus/tree/socs24-width-and-mq. Following (chen-thiebaux:socs-2024) we compare against $^{\text{md}}, h+h$$^{\text{mrp}}_{\text{+hj}}$, $h^{\text{add}}$, their novelty variants $^{\text{md}}_{\langle\text{B, QB}\rangle}, h+h$$^{\text{mrp}}_{\text{+hj}~\langle\text{B, QB}\rangle}$, $h^{\text{add}}_{\langle\text{B, QB}\rangle}$, as well as the three portfolio settings, $P(3h)$ (non-novelty only), $P(3n)$ (novelty only), and $P(3h\,\|\,3n)$ (both). We report of them only $h^{\text{md}}$, the best-performing single heuristic $h^{\text{add}}_{\langle\text{B, QB}\rangle}$, and the best portfolio configuration, $P(3h\,\|\,3n)$, while the full comparison is provided in Appendix.
Metric-FF (MFF) (hoffman:jair-2003) is used off-the-shelf. Implemented in C, it employs Enforced Hill Climbing Search (EHCS) with the interval-relaxed $h_{\text{FF}}$ heuristic and Helpful Operators.777fai.cs.uni-saarland.de/hoffmann/metric-ff.html The planner is incomplete and may report problems as unsolvable due to limitations of EHCS.
Numeric-FD (NFD) (aldinger-nebel:socs-2017) is a C++ planner that employs Lazy GBFS with numeric LM-cut heuristics (kuroiwa-et-al:icaps-2022).888github.com/ipc2023-numeric/team-1 This planner won the numeric track of the IPC-2023 (see taitler-etal:aimag-2024).
The Pacman and Twin Prime domains demand a level of expressiveness that makes encoding them for existing planners highly impractical and cumbersome in practice. As there currently are no available methods to derive a domain-independent heuristic for this type of tasks, we benchmarked them against BFS and $h^{\text{md}}$.
LLM-generated heuristics are not always compilable and vary in quality and efficiency generation-to-generation. To address this, we employ three fallback strategies:
In FirstCompilation (fc), we repeatedly query the LLM until a compilable heuristic is obtained and run GBFS with it until a solution is found or resources are exhausted, 600 sec (search + all API calls).
In UntilSuccess (us), we operate the same as fc, except we continue even if a compilable heuristic failed.
In SelfPortfolio-N (sp-n), we allocate fixed $Time/N$s slices to N heuristics; if a run fails or exceeds memory, we restart with generating a new heuristic. This continues until a solution is found or all slices failed. All API call durations, typically 10–45 seconds depending on the model, are included in the time budget.
We evaluated The following LLMs and settings: OpenAI’s GPT-4.1 and GPT-4.1-mini, their reasoning models GPT-5.1 and GPT-5-mini (under both a ”low” and ”high” reasoning effort setting), and Anthropic’s Claude Sonnet 4.5 and Claude Haiku 4.5. Results are presented in Table 1. For practicality, we limit the maximum number of heuristics generated for fc and us to 10, which if passed the instance is declared failed. As is standard in AI planning, we report coverage rather than averages, since state space size typically grows exponentially and averaging would be misleading. As a qualitative measure of the generated heuristics we report the head-to-head time and expanded nodes for solution comparison of sp-10 with GPT-4.1 vs. the best performing $h^{\text{add}}_{\langle\text{B, QB}\rangle}$, which can be seen in Figure 5.Comparisons with additional models can be seen in Appendix.
We conducted ablation studies to evaluate the contribution of three components in our heuristic generation pipeline:(i) the provision of the specific instance (ii) the model and fallback choice (iii) the reasoning effort requested from reasoning models. The study aimed to determine their impact on heuristic quality, generation time, and planning performance, as well as the variance in our best configuration.
Impact of Instance-Specificity
We measure the effect of providing the instance JSON to the model before generation, containing the initial state and the goal. This comes at the steep cost of requires generating additional heuristics per instance, as opposed to domain-specific-instance-general which can be generated once then ran on all future instances. Intuitively, IS may have the following benefits:
  1. 1.
    The heuristic often contains multiple penalizing elements which are combined with guessed coefficients. Knowing the details of the instance gives the model a better idea on how to tune these coefficients.
  2. 2.
    It tells the model how significant each part of the problem is. Consider a simple instance of Pacman where the ghosts are trapped. The model can avoid spending generation- and run-time on a complex subroutine to quantify the risk from the ghosts.
  3. 3.
    It tells the model the scale of the problem which can provide guidance on efficiency constraints. We hypothesize that there exists a tradeoff between the precision of the heuristic estimates and the computation time, and further that for larger problems it can be worthwhile for the heuristic to sacrifice precision for efficiency. For example, on TPP, Anthropic’s models implement path-finding algorithms like Dijkstra, Prim and Floyd-Warshall, about half the time. Although this is a good basis for a “smart” heuristic, it causes a timeout within a few thousand expansions for even medium problems. Meanwhile even Manhattan Distance heuristics achieves 16/20.
  4. 4.
    At times, a solution state is obvious from the initial state (e.g, FO-Counters). In these cases showing the model the instance details allows it to write a heuristic to guide efficiently toward that solution rather than abstractly guiding toward all solutions.
Empirically, IS leads to a decrease in overall coverage (Table 2) with most domains varying slightly but a few increasing or decreasing significantly. IS did not cause heuristic generation time to increase [^9] Interestingly, despite adding tokens and theoretically adding complexity which can increase output tokens, adding the instance slightly decreased generation time. (P=2e-5<<0.05). We do not have an explanation for this fact.. IS moderately increased compilation errors (see Table  3), and if this effect is normalized against, IS-heuristics have a slightly higher solving rate. We hypothesize that GPT-4.1 is insufficiently strong to make use of this info and that for a stronger model it may be worth revisiting.
FC US SP-10
Domain $D$ IS $D$ IS $D$ IS
Block Grouping (20) 16 9 19 10 19 10
Counters (20) 10 8 10 11 10 11
Delivery (20) 16 10 20 18 20 18
Drone (20) 15 15 16 16 16 16
Expedition (20) 3 1 3 4 3 4
Farming (20) 18 19 19 20 19 20
FO-Counters (20) 4 8 4 14 4 14
FO-Farming (20) 20 18 20 20 20 20
FO-Sailing (20) 20 10 20 19 20 20
Hydropower (20) 8 7 12 9 12 9
Market Trader (20) 20 1 20 3 20 3
Pathways (20) 0 1 1 1 1 1
Plant Watering (20) 9 12 20 18 20 18
Rover (20) 3 2 4 3 4 3
Sailing (20) 11 14 20 20 20 20
Settlers (20) 1 2 1 3 3 3
TPP (20) 13 11 16 13 16 15
Zenotravel (20) 20 14 20 20 20 20
$\sum$ (360) 207 162 245 222 247 225
Twin Prime (20) 18 20 18 20 18 20
Pacman (20) 9 13 9 13 16 14
$\sum$ (400) 234 195 272 255 281 259
Table 2: Coverage results for GPT-4.1 with domain-dependence only ($D$), our standard configuration, against GPT-4.1 with Instance-Specificity (IS), reporting each on fc, us, and their best configuration, SP-10. For most domains GPT-4.1 fails to leverage the additional information, however we highlight FO-Counters and Twin Prime as domains where it lead to the most improvement.
Impact of model and fallback choice
We tested 6 models: GPT-4.1, GPT-4.1-mini, GPT-5.1, GPT-5-mini, Claude Sonnet 4.5, and Claude Haiku 4.5, each with each of our fallback options including fc, us, sp-5, and sp-10. For the full table see Appendix. Our observations are as follows:
  1. 1.
    For non-reasoning models, under about every configuration, a model’s larger variant is preferred. Since they are well-capable of generating and running many heuristics within the allotted time, SP10SP-10 is preferred.
  2. 2.
    Reasoning models are more competitive between sizes. They take longer to generate and $US$ performs best, allowing them the time they need.
  3. 3.
    Anecdotally, when reasoning models fail it can often be because they over-complicate the problem, taking longer to generate, and whether they end up with a better heuristic or an unusable one is a coin toss. That said, since we can re-sample this risk may be worth taking if the samplings are sufficiently independent.
Impact of Reasoning effort
We tested our method on the same model, GPT-5.1 with two configurations: reasoning_effort=’low’ and reasoning_effort=’high’ (for brevity denoted ’GPT-5.1’ and ’GPT-5.1=high’). We found that although increasing reasoning_effortreasoning\_effort slightly increased heuristic efficiency and quality, and better-than-halved compilation error rate, when we account for generation time the reduced number of heuristics we generate in time does not make this trade worthwhile. With better engineering one could generate many completions simultaneously, which may turn this trade worthwhile. We leave this direction for future work. For further graphs see appendix.
GPT Claude
Metric 4.1 4.1 (IS) 4.1 mini 5.1 5 mini 5.1 high 5 m-h Sonnet 4.5 Haiku 4.5
Input tok. 3725 8102 3725 3724 3724 3724 3724 4761 4761
Output tok. 1510 1695 1245 3253 2298 18075 14485 2122 1490
Reasoning tok. [^10] For GPT, subset of output tokens. For Claude, they are not provided by the API - - - 1282 589 15510 11613 - -
Cost (USD) 0.02 0.03 0.003 0.037 0.006 0.185 0.03 0.046 0.012
Generation (sec.) 35.8 27.5 19.4 44.9 44.5 235.1 202.3 33.6 12.5
Run (sec.) 24.7 22.1 19.2 24.2 18.0 21.4 36.7 43.4 22.2
Compl. error % 0.300 0.430 0.215 0.150 0.405 0.070 0.205 0.195 0.165
OOM rate % 0.304 0.241 0.388 0.382 0.280 0.340 0.288 0.391 0.451
Runtime error % 0.027 0.020 0.035 0.032 0.028 0.035 0.032 0.016 0.032
Timeout % 0.015 0.014 0.004 0.007 0.009 0.007 0.031 0.032 0.002
Success % 0.355 0.295 0.358 0.430 0.278 0.548 0.444 0.365 0.351
Table 3: Average token usage, cost, latency, and execution outcomes for different GPT and Claude models, independent of downstream applications. Models marked with are reasoning models. Excepting IS, input tokens across models depend only on the tokenizer. Costs do not account for token caching discounts.
Figure 4: As a variance analysis we generated 40 heuristics with GPT-4.1 in each domain, and performed a Monte-Carlo simulation of our algorithm for 1000 iterations of sampling heuristic order, taking the coverage per-domain and overall each time. The boxes represent the first and third quartiles, the thick line the median, and the whiskers the 1st and 99th percentiles. Although limited, the analysis shows the method to be highly consistent for most domains.
The ablation results suggest that non-reasoning models don’t significantly benefit from IS. The best setting of our method is to use a non-reasoning model to generate non-IS heuristics and run them as SelfPortfolio-10 (276 with GPT-4.1). If time is allowed ahead of receiving the instances, the best setting becomes instead to generate ahead of time as many heuristics as possible and run SP-N or US on them (The two fallbacks are approximately equivalent when used on pregenrated heuristics due to near-zero timeout rate).
Discussion
As shown in Table 1, LLM-based heuristics achieve state-of-the-art performance across the evaluated domains. They are particularly effective in challenging benchmarks such as (FO-) Sailing and Zenotravel, which are known to be difficult for domain-independent planners. The best-performing LLM-based heuristics were produced by GPT-5.1 and it’s mini variant with high reasoning effort, however when accounting for generation time and cost, GPT-4.1 emerges as a clear winner. Nonetheless, certain domains remain challenging for LLMs, likely due to their sensitivity to non-descriptive variable names and opaque goal specifications, as seen in Pathways and Settlers.
Expressiveness
In the Pacman and Twin Prime domains, no available planner can solve problems requiring this level of expressive power, necessitating a comparison against BFS and $h^{\text{md}}$. As expected, LLM-based heuristics outperformed both. Overall, LLM-based heuristics surpass state-of-the-art numerical planners while addressing problems that cannot be adequately expressed in any existing pddl dialect.
Representation
One concern is that the performance gains observed with LLM-generated heuristics may result from optimized planner code rather than the heuristics themselves. To test this, we reimplemented $h^{\text{md}}$ within our planner for a direct comparison with ENHSP-20 using the same heuristic. As shown in Table 1, ENHSP-20 with $h^{\text{md}}$ generally achieves better results, particularly in domains such as Hydropower and Plant Watering, which are highly sensitive to problem representation. This suggests that departing from pddl-based input to a direct successor-generator representation may, in some cases, degrade performance, allowing us to reject the hypothesis that our representation alone accounts for the observed improvements.
Costs and Efficiency While concerns about the resource efficiency of LLMs have been raised (stojkovic2024towards), energy efficiency is beyond the scope of this work. Nonetheless, our method is amenable to efficient deployment: generating heuristics can be done ahead of time one per domain then reused on every instance, substantially reducing energy and cost even if a powerful model is used. Additionally, the resulting heuristics tend to be very efficient to run, and succeed or fail much quicker than domain-independent heuristics (see Appendix.)
Conclusion
This paper explored the use of LLMs to generate heuristic functions directly from AI planning task definitions, bypassing traditional domain-independent heuristics written in formal languages like pddl. Our approach uses general-purpose code (Rust) to represent planning tasks and lets LLMs synthesize tailored heuristics.
Empirically, the results show that LLM-generated heuristics achieve state-of-the-art performance across several established planning benchmarks. The approach excels in domains traditionally difficult for domain-independent planners, such as Zenotravel, indicating substantial potential for efficiency gains. Furthermore, it successfully addresses complex planning tasks that are not easily expressible in conventional formalisms, like Pacman and Twin Prime.
Nonetheless, limitations remain. LLM-generated heuristics are sensitive to the clarity and descriptiveness of task representations; unintuitive state-space encodings degrade their quality. Moreover, while larger models such as GPT-4.1 and Sonnet 3.5 tend to produce high-quality heuristics, they incur greater computational costs. Smaller models offer improved cost-efficiency but often generate less reliable code. This trade-off highlights the importance of selecting models according to task complexity and scalability requirements.
Looking forward, this work opens several promising avenues for future research. These include improving robustness across domains, comparing LLM-generated heuristics to hand-crafted ones, and exploring hybrid strategies combining LLMs with domain knowledge or iterative refinement techniques.